OpenStudy (anonymous):

how to find the value of cot 30 without using tables / calculator?

OpenStudy (anonymous):

cot(30) Take the cotangent of 30 to get 1 3‾√ . 1 3‾√ To rationalize the denominator of a fraction, rewrite the fraction so that the new fraction has the same value as the original and has a rational denominator. The factor to multiply by should be an expression that will eliminate the radical in the denominator. In this case, the expression that will eliminate the radical in the denominator is (3)‾‾‾√ (3)‾‾‾√ . 1 3‾√ ⋅ (3)‾‾‾√ (3)‾‾‾√ To eliminate the radical from the denominator, multiply 3‾√ by 3‾√ to get 3. (3)‾‾‾√ 3 Simplify the rationalized expression. 3‾√ 3 this is simplified if thats what your asking

OpenStudy (anonymous):

simplified is equal to the value

OpenStudy (y1066380):

cot30 =1/tan30 =1/(sqrrt(3)/3) =3/sqrrt(3)